OpenStudy (anonymous):

Please Help with Algebra 1 Thanks MEDAL

OpenStudy (anonymous):

Which equation is equivalent to y – 6 = –12(x + 4)? (1 point) y = –6x – 48 y = 6x – 48 y = –12x – 42 y = –12x – 54 2. Which point is located on the line represented by the equation y + 4 = –5(x – 3)? (1 point) (–4, –5) (–5, –4) (3, –4) (–3, 4) 3. Which equation represents the line that passes through the points (6, –3) and (–4, –9)? (1 point) y + 4 = (x + 9) y + 4 = (x + 9) y - 3 = (x + 6) y + 3 = (x – 6) 4. The population of a city increases by 4,000 people each year. In 2025, the population is projected to be 450,000 people. What is an equation that gives the city’s population p (in thousands of people) x years after 2010? (1 point) p = 4x + 50 p – 450 = 4(x – 15) p – 15 = 4(x – 450) p = 4x + 15 5. Graph the equation. y – 3 = –(x + 4) (1 point)

OpenStudy (anonymous):

i like cheez

OpenStudy (anonymous):

Thanks for that do you go to Connections Academy?